OpenStudy (anonymous):

Please help! Medal/fan Simplify the expression (x^1/3)^2/(x^3)^8/3

jhonyy9 (jhonyy9):

(x^1/3)^2 = (x^2/3) (x^3)^8/3 = (x^24/3) = x^8 (x^1/3)^2 (x^2/3) ---------- = -------- = x^((2/3) -8) = x^((2-24)/3) = (x^3)^8/3 x^8 1 1 = x^((-22)/3) = --------- = -------------- (x^22/3) (x^7)*(x^1/3)